OpenStudy (anonymous):

see note A) (-∞, ∞) B) [5, ∞) C) (-∞, 5) (5, ∞) D) (5, ∞)

OpenStudy (anonymous):

x\[\frac{ x }{ \sqrt{x-5} }\]

OpenStudy (anonymous):

x-5=0 x=5 so C

OpenStudy (anonymous):

RIGHT!

OpenStudy (anonymous):

YES? NO?